OpenStudy (anonymous):

(square soot of -2)(square root of 2)+(square root of -4i)- 2i^6+(square root of -2i^2)

OpenStudy (anonymous):

\[\sqrt{-2}\sqrt{2}+\sqrt{-4i}-2i ^{6}+\sqrt{-2i ^{2}}\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you just supposed to simplify

OpenStudy (anonymous):

yes, the sqare roots of -2 and 2 i dnt know how to do

OpenStudy (anonymous):

Ok. I'll give it a launching off point

OpenStudy (anonymous):

\[\sqrt{(-1)(2)}\rightarrow \sqrt{-1}\sqrt{2}\rightarrow i \sqrt{2}\rightarrow so \rightarrow i \sqrt{2}\sqrt{2}\]

OpenStudy (anonymous):

can you see that the you will then end up with 2i

OpenStudy (anonymous):

thts for both sqaure roots?

OpenStudy (anonymous):

the first part is for the negative. After the so I included the positive root. I was just doing one at a time. Thats only for the first term. Do you know how to do the remaining terms?

OpenStudy (anonymous):

i see...

OpenStudy (anonymous):

can you wait a sec, please

OpenStudy (anonymous):

sqaure root of -4i is \[i \sqrt{2i}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[2i \sqrt{i}\]

OpenStudy (anonymous):

-2i^=2.. so that +2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

+ \[i \sqrt{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do \[i ^{2}\]first, try again.

OpenStudy (anonymous):

+\[\sqrt{2}\]?

OpenStudy (anonymous):

yup. I don't think you can simplify it anymore from here. good job. I just ask if you could give me a medal. Makes me more reputable.

OpenStudy (anonymous):

sure, thanks alot

OpenStudy (anonymous):

where do i go to?

OpenStudy (anonymous):

:)